OpenStudy (anonymous):

Lucas recorded his lunch expenditure each day for one week in the table below. Day: Sunday ($4.85), Monday($5.10), Tuesday($5.50), Wednesday ($4.75), Thursday ($4.50), Friday($5.00) Saturday($6.00) a. calculate the standard deviation using the formula Standard-Deviation.png (1063×598) round to the nearest thousandth. b. If the sample size increases but the sum of the squared differences stays the same, what will be the effect on the standard deviation?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

@godzgyrl2001

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@katlin95 @geerky42 @blurbendy @e.mccormick

OpenStudy (anonymous):

@Peter14

OpenStudy (e.mccormick):

Do you understand the formula?

OpenStudy (anonymous):

Not at all. I am completely lost.

OpenStudy (e.mccormick):

With explanaitory notes: http://0.tqn.com/d/statistics/1/0/M/-/-/-/standarddev.GIF

OpenStudy (e.mccormick):

Does that help?

OpenStudy (anonymous):

I get the first one but not the other two.

OpenStudy (e.mccormick):

Sigma, \(\Sigma\), means sum.

OpenStudy (e.mccormick):

OK. Let's jsut walk through it together. So you need to start by finding the mean. Know how to do that?

OpenStudy (anonymous):

Yeah, it's just the average right?

OpenStudy (e.mccormick):

Yes. That is the arithmetic mean.

OpenStudy (anonymous):

I got 5.1

OpenStudy (e.mccormick):

In this formula that is \(\overline{x}\). I got 5.1 too. So we will remember that for now.

OpenStudy (e.mccormick):

Next, we need the deviation of each value from the mean, that means for each item, or \(x_i\), we subtract 5.1 and write those new values down. The devations need to be squared, then added. That is the whole: \(\sum_1^n(x_i-\overline{x}})^2\) part.

OpenStudy (anonymous):

okay, what next?

OpenStudy (anonymous):

Okay that just confused me

OpenStudy (e.mccormick):

Man that was a laggy bit.... OK, so. lets just start with the deveation part. You have: Sunday ($4.85), Monday($5.10), Tuesday($5.50), Wednesday ($4.75), Thursday ($4.50), Friday($5.00) Saturday($6.00) For each one, what is that value \(-5.1\)

OpenStudy (anonymous):

-.25 0 .4 -.35 -0.6 -.1 .9?

OpenStudy (e.mccormick):

Devation means, "How far away from the mean." That is why we are subtracting the mean. Yes. Those are the deveations. Now, square each one of them.

OpenStudy (e.mccormick):

Then add up those values. The squared ones. 0.0625+ 0+ 0.16+ 0.1225+ 0.36+ 0.01+ 0.81= 1.525 I think you know basic math, so there is that one for you.

OpenStudy (e.mccormick):

That entire process was: Find the mean. Find each deviation from the mean. Square each deveation. Sum the squares. In math, that is done in shorthand like this:\[\sum_1^n(x_i-\overline{x})^2\]All that says is do what we just did. OK? That is probably the most complex part of this all. There is still some left, but that was the big part.

OpenStudy (e.mccormick):

The site has been bumping people off for a while, and it looks like you got zapped. OK, we started with this formula:\[S_x=\sqrt{\frac{\sum_1^n(x_i-\overline{x})^2}{n-1}}\] Do far we found that \(\sum_1^n(x_i-\overline{x})^2=1.525\). So if we put that into the formula it becomes:\[S_x=\sqrt{\frac{1.525}{n-1}}\] n is the number of entries, so 7 in this case. That should be enough info for you to finish it. You got the hard part before getting zapped.

OpenStudy (anonymous):

my brain is processing sorry lol.

OpenStudy (e.mccormick):

You made it back! Wheee.

OpenStudy (e.mccormick):

Give that all a read over, then see if you get it or if you have questions.

OpenStudy (anonymous):

thank you so much!!!!!!!!

OpenStudy (e.mccormick):

I take it that means it made sense. Like I said, the sigma notation sum is the conceptually hard part. The math of a sigma sum is not hard. We did it. But until you know the concept, how can you do the math?

OpenStudy (e.mccormick):

Another thing to notice is that the absolute value of each of the devations is not too far from the answer. The answer is what is called the Standard Deveation. When you realize that the Standard Deveation is made of the individual deveations, the name makes mroe sense.

OpenStudy (anonymous):

So what about the second part?

OpenStudy (e.mccormick):

Well, you see the [\S_x=\sqrt{\frac{1.525}{n-1}}\]Can you solve that now?

OpenStudy (e.mccormick):

oops. It neded to start that with slash brace. hehe. \[S_x=\sqrt{\frac{1.525}{n-1}}\]

OpenStudy (anonymous):

I got part a but I'm not sure how to do part b.

OpenStudy (e.mccormick):

Well, n is the number of days, so 7. Put that in for n. After that, it is just math.

OpenStudy (anonymous):

I don't understand...

OpenStudy (anonymous):

Just to clarify, the answer to part a is 0.46675 right?

OpenStudy (e.mccormick):

OH! That part b... ooops.... I had not scrolled back all the way up.... I got 0.5041494487 for a.

OpenStudy (anonymous):

Hmmmmmmm. how?

OpenStudy (anonymous):

Oh wait! the formula I have on the bottom has n not n-1

OpenStudy (e.mccormick):

\[S_x=\sqrt{\frac{1.525}{n-1}}\implies S_x=\sqrt{\frac{1.525}{7-1}} \implies S_x=\sqrt{\frac{1.525}{6}} \implies \\ \implies S_x=\sqrt{0.2541666667} \implies S_x=0.5041494487\]

OpenStudy (e.mccormick):

OH. OK. The difference is for globals vs. locals. They aare both SD formulas. the one you linked in the picture is the n-1 version, which would be a local SD.

OpenStudy (anonymous):

So I'm actually looking for the global SD?

OpenStudy (e.mccormick):

It does not matter what it is called, as long as you use the right formula. If that formula you linked is what they designated, you MUST use n-1.

OpenStudy (anonymous):

my worksheet has n on bottom

OpenStudy (anonymous):

I got 0.5 for the answer.

OpenStudy (e.mccormick):

OK, with n, I got 0.4667516929, which in thousanths is .467

OpenStudy (e.mccormick):

Now, the sample size is n. So if n is larger... we did n-1, which would be smaller. So n being larger is 8....

OpenStudy (anonymous):

okay now what about the 2nd part?

OpenStudy (e.mccormick):

Try it with like n=8 or n=10. See what happens.

OpenStudy (anonymous):

but how does the sample size get bigger and the sum of the squared differences stay the same ?

OpenStudy (e.mccormick):

it is a theorhetical question.

OpenStudy (anonymous):

would the SD stay the same?

OpenStudy (e.mccormick):

It is a "what if" they pulled out of their anatomy in an effort to get the students to see something... even if the student is smart enough to say, "But that would change thing, so it is a bad question!" they still want it ansewered. No, it would chamnge. Just change n and see.

OpenStudy (anonymous):

I don't know how to add more numbers but keep the sum the same.

OpenStudy (e.mccormick):

No, no. Just change it like this: \[S_x=\sqrt{\frac{1.525}{10}}\]

OpenStudy (e.mccormick):

The problem is you were too smart for the question! It is a stupid question and it does not make sense. But that is what they want you to do.

OpenStudy (anonymous):

Oh! haha thank you! so it should go down right?

OpenStudy (anonymous):

cause I got 0.391

OpenStudy (e.mccormick):

yah. SD is gets smaller when that happens. Also, SD gets more accurate when the number gets larger. That second concept is something they are probably leading up to, but at this point they want you to see that it gets smaller.

OpenStudy (anonymous):

Thank you so much! I will probably tag you in more tomorrow!

OpenStudy (e.mccormick):

If the site works... this has been.... ugly tonight.

OpenStudy (anonymous):

Definitely. One more thing, it asks for a real world application?

OpenStudy (e.mccormick):

Wheee... got logged out again. Well, find a range of money needed for a lunch budget... hmmm.... Stats is more along the lines of finding how one person or a small group in comparison to normal.

OpenStudy (anonymous):

Alright, thanks for all your help!

OpenStudy (e.mccormick):

np. Have fun!